# Training objective drives the consistency of representational similarity across datasets

### v1

@@ Appendix F Validating training objective effects through controlled anchor model analysis
Appendix G Distribution correlation coefficients downstream task performance vs. model similarity
@@
In Sec. 4.7, we examined the correlations between representational similarity and classification performance gaps across three datasets from each of the four dataset categories: Natural (multi-domain), Natural (single-domain), Specialized, and Structured. We observed stronger negative correlations for the Natural (single-domain) datasets. Fig. 15 extends this analysis by illustrating the correlation coefficients for all evaluated datasets. We can confirm that natural (single-domain) datasets exhibit the strongest negative correlations overall on this larger set of datasets, with the SVHN dataset being the sole outlier.
@@
While our categorization of natural datasets into single-domain and multi-domain revealed an interesting difference in the relationship between downstream task performance and model similarity—especially between the ImageNet-1k and single-domain datasets like Flowers or Pets—the correlation coefficients for these categories still overlap. We hypothesize that our binary sub-categorization of Natural datasets is insufficient for capturing a clear distinction for two reasons: (1) Datasets lie on a spectrum between the two (hypothetical) extremes of encompassing all possible domains and encompassing exactly one. For instance, ImageNet-1k has a relatively rich set of domains, containing 1.000 diverse classes, whereas CIFAR-10 only contains 10 classes of vehicles and animals. (2) Datasets differ in the amount of contextual information (e.g., the object’s background) they offer. We assume that contextual information plays an important role in the number of viable features a dataset classification can be solved with. A large number of viable features means that multiple well-performing but dissimilar strategies may exist. Consequently, multi-domain datasets with limited contextual information may behave more like single-domain datasets. For example, the PASCAL VOC 2007 dataset, when used for single-label classification, makes use of tight bounding-box crops around the object to be classified, reducing the available context. Similarly, in low-resolution datasets like CIFAR-10 and CIFAR-100, contextual information may be reduced due to the loss of background information. While we have not quantified it, we found the Caltech-101 dataset to contain much cleaner object backgrounds than ImageNet-1k, which also suggests reduced contextual information.
Figure 15: Correlation coefficients between downstream task performance and model representational similarity (CKA linear) grouped by dataset categories measured with Pearson correlation.

### v2

@@ Appendix I Effect of training data domain on similarity consistency
Appendix J Distribution correlation coefficients downstream task performance vs. model similarity
@@
In §4.7, we examined the correlations between representational similarity and classification performance gaps across three datasets from each of the four dataset categories: Natural (multi-domain), Natural (single-domain), Specialized, and Structured. We observed stronger negative correlations for the Natural (single-domain) datasets. Fig. 21 extends this analysis by illustrating the correlation coefficients for all evaluated datasets. We can confirm that natural (single-domain) datasets exhibit the strongest negative correlations overall on this larger set of datasets, with the SVHN dataset being the sole outlier.
@@
While our categorization of natural datasets into single-domain and multi-domain revealed an interesting difference in the relationship between downstream task performance and model similarity—especially between the ImageNet-1k and single-domain datasets like Flowers or Pets—the correlation coefficients for these categories still overlap. We hypothesize that our binary sub-categorization of Natural datasets is insufficient for capturing a clear distinction for two reasons: (1) Datasets lie on a spectrum between the two (hypothetical) extremes of encompassing all possible domains and encompassing exactly one. For instance, ImageNet-1k has a relatively rich set of domains, containing 1,000 diverse classes, whereas CIFAR-10 only contains 10 classes of vehicles and animals. (2) Datasets differ in the amount of contextual information (e.g., the object’s background) they offer. We assume that contextual information plays an important role in the number of viable features a dataset classification can be solved with. A large number of viable features means that multiple well-performing but dissimilar strategies may exist. Consequently, multi-domain datasets with limited contextual information may behave more like single-domain datasets. For example, the PASCAL VOC 2007 dataset, when used for single-label classification, makes use of tight bounding-box crops around the object to be classified, thereby reducing the available context. Similarly, in low-resolution datasets such as CIFAR-10 and CIFAR-100, contextual information may be reduced due to the loss of background information. While we have not quantified it, we found the Caltech-101 dataset to contain much cleaner object backgrounds than ImageNet-1k, also suggesting reduced contextual information.
Figure 20: Model similarity (CKA linear) vs. absolute difference in downstream task performance (top-1 accuracy) for model pairs for three datasets per dataset category ( natural multi-domain, natural single-domain, specialized, and structured). The color of each point indicates the downstream task accuracy of the better-performing model. Pearson correlation coefficients are shown in each subplot.
Figure 21: Pearson correlation coefficients between downstream task performance gap and model representation similarity for all 23 datasets grouped by dataset category.